OpenStudy (anonymous):

could anybody help me with this.. i have know idea what to do. Solve the equation: -4(x + 10) - 6 = -3(x - 2)

OpenStudy (anonymous):

\[-4x-40-6=-3x+6\] \[-4x+3x=6+40+6\] \[-x=52\] \[x=-52\]

OpenStudy (anonymous):

a ( b + c ) = ab+ac

OpenStudy (anonymous):

-a ( b + c ) = -ab-ac

OpenStudy (anonymous):

oh it makes sense thanks.